A room at 20 degree celsius is heated by heater of resistance 20 ohm connected to 200 volt mains. The temperature is uniform and it is transmitted by a glass window of area 1 m square and thickness 0.2 cm. Calculate the outside temperature? (Kglass=0.2/•c)

Grade:11

4 Answers

souradeep
6 years ago
hhh h h hh hhhhh fg fggf  gf f g fg fg fg fg fg fg fg fgf g gf gfggfmgf gffgfgfg gffg fggf gf gf gf gf g gf fg gf gf gffg gf gf gf gf gf g  gff gffg
souradeep
6 years ago
nnmmn   vhggyu yugk   kug yg guy  kguy kgyu kguy ykug kgyu yku y gkuy gku yygu k g yku ykuggku y yku yu y  uyk
P
6 years ago
Q/t=i=V/R=200/20=10 amperesQ/t=K×area×temperaturedifference/thicknessi=0.2per degree celsuis×1m×T-20degree celsius/0.2cm10=100(T-20)T=20.1
souradeep
6 years ago
n nn nn n n n nn n n nn n  n nn n n n nn n n kjkvckh kkjm kjm kh hn n hn hn khn khn khn khn khn khn hn  hh  hhhnk yfvdrv dvytv ytbf nki